OpenStudy (anonymous):

Someone please help me! How do you simplify this rational expression? 4a+4/a+3 * 3a+9/2a-6

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

\[\frac{ 4a+4 }{ a+3 } \times \frac{ 3a+9 }{ 2a-6 }\]